OpenStudy (anonymous):

What is the graph of the system y = –2x + 3 and 2x + 4y = –8?

OpenStudy (anonymous):

y = –2x + 3 m=-2 y-int (0,3) 2x + 4y = –8 ===> y = -1/2x - 2 m=-1/2 y-int(0,-2) can you finish it now?

OpenStudy (anonymous):

yea thank you!